OpenStudy (anonymous):

solve for x 2x-4 /10 = 2x / 20 Will fan

OpenStudy (calculusxy):

\[\frac{ 2x - 4 }{ 10 } = \frac{ 2x }{ 20 }\]

OpenStudy (calculusxy):

Is that the equation?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

x=4

OpenStudy (calculusxy):

So we can see that this is a proportion.

OpenStudy (calculusxy):

Both of the fractions are equal to each other. And knowing that this is a proportion will help us.

OpenStudy (calculusxy):